# Anonymization by Design of Language Modeling

### v1

@@ 4. Privacy-Preserving Language Modeling
In order to prevent the model from memorizing personal information, the privacy preserving language modeling that we propose ensures that the choice of mask used during specialization avoids directly and indirectly identifying words. To achieve that, this solution is based on a preprocessing that is responsible for identifying both the direct and indirect identifying words. We first present this preprocessing step (Section 4.1) before explaining how the masked and the causal language modeling work (Section 4.2 and Section 4.3, respectively).
@@ 4.1. Preprocessing: Building a Blacklist
The preprocessing step focuses on identifying identifiers that the model must not memorize. First, we use an off-the-shelf NER to directly identify identifying words associated with certain categories. This NER can be adapted to include more or less categories depending on the field of specialization. Then, we build the bipartite graph $G$ between the individuals of the corpus and the words used in their documents in order to identify indirectly identifying words (similar to Figure 3). To do this, once the graph is built, we identify the words that are pointed to only by one individual. Finally, all these (directly or indirectly) identifying terms are kept in a kind of blacklist that is used when modeling the masked or causal language.
If we make the link with the notion of k-anonymity(Sweeney, 2002), which defines that a data is anonymous if it is used by at least k different users, all the words used as masks to specialize the model in the language modeling step are at least used by two people (i.e., k=2). This makes the task of an adversary wishing to infer the membership of an individual more difficult. We can easily change the reading of the bipartite graph to consider a larger value of k for the identification of indirectly identifying information (e.g., identifying words that are only pointed to by 1 or 2 users to have a k=3).
In addition to being able to vary the value of k to protect indirectly identifying words and to be able to use a NER adapted to the domain of specialization to identify directly identifying words, it should be noted that the cost of this preprocessing is reduced.
@@ 4.2. Privacy-Preserving Masked Language Modeling
Compared to standard masked language modeling that randomly chooses terms to mask during supervised learning, our privacy-preserving masked language modeling ensures that directly or indirectly identifying terms are never masked by leveraging the blacklist constructed during the preprocessing step. More precisely, the choice of the mask is made randomly among the words that are not in the blacklist.
@@ 4.3. Privacy-Preserving Causal Language Modeling
Similar to privacy-preserving masked language modeling, our privacy preserving causal language modeling ensures that masked words are not directly or indirectly identifying words. However, since masked words are always at the end of the sequence to be predicted, we exploit the blacklist created during the preprocessing step to check that this last term is not an identifier. If it is, we modify it and replace it with a padding token and add the next word for prediction. The advantage of using a padding token is that it is not used in the loss calculation during training, and therefore does not degrade the model’s predictions.
@@ 5. Experimental Setup
We performed an extensive evaluation of our privacy-preserving language modeling scheme using medical datasets (Section 5.1) and considering both BERT and GPT based language models (Section 5.2). To capture the full impact of our solutions, we considered a set of both utility and privacy metrics (Section 5.3) and considered different baselines approaches (Section 5.4). Finally, we describe our methodology (Section 5.5).
@@ 5.1. Datasets
To conduct our experiments, we considered medical field datasets, specifically the N2C2 (National NLP Clinical Challenges) datasets (Henry et al., 2020) and a dataset from the Hospices Civils de Lyon (HCL) (Richard et al., 2023).
  • N2c2: we leverage two datasets gathering medical discharge summaries (i.e., english free text). In the first one (Uzuner et al., 2007), 928 records have been annotated for replacing all authentic Personally Identifiable Information (PII) with realistic surrogates. These PIIs fall into different categories: six of the seventeen textual PII categories listed by HIPAA (only these six categories appear in the data: patients, locations, dates, IDs, phone numbers, and ages), and two additional categories, doctors and hospitals, resulting in eight PII categories in the dataset. In the second dataset (Uzuner et al., 2008) almost 500 records have been annotated by pulmonologists. The pulmonologists were asked to classify patient records into five possible smoking status categories as past smoker, current smoker, smoker, non-smoker, and unknown. We leverage this dataset to assess the utility degradation of our privacy-preserving language modeling through a downstream classification task.
  • Hcl: this dataset gathers french free text from medical reports, notes, prescriptions, etc. This dataset contains texts from 1240 patients hand-annotated by ten people, with double verification to identify PII falling into the following categories: Name, First name, Dates, Organizations, Telephones, Emails, Cities, Postal Codes, Streets, File Number, Website, Localities and IPP.
@@ 5.2. Model Architecture
In order to cover the different use cases of language models (i.e., the generative models and the descriptive text models), we consider both the BERT architecture (i.e., the base model case composed of 12 layers, 768 hidden dimensions, 12 attention heads, 110M parameters) and the GPT-2 architecture (Radford et al., 2019) (the smallest version of GPT-2, with 124M parameters) from Hugging Face222 Hugging Face - https://huggingface.co/. To support the French dataset, we also considered a CamemBERT (i.e., a BERT model specialized for French) and a GPT model specialized for French. We did not consider larger model versions due to resource limitations, and we also discarded distilled version of these models (Sanh et al., 2020b).
@@ 5.3. Metrics
We consider both utility and privacy measures in our evaluation to better assess the trade-off between protecting the training data and the model’s performance in specializing on a specific corpus. In particular, for the utility evaluation, we quantify the ability of the models to predict masked terms well. For the utility evaluation of BERT-like models, we also quantify the model’s performance in classification by creating a downstream task. For GPT-like models, the utility is evaluated by the model’s ability to generate the right tokens in the right places. To evaluate the privacy of the models, we consider their ability to predict identifiers (i.e., direct and indirect) contained in the dataset used to specialize the model. Since a privacy leakage is marked when an identifier is predicted by the model regardless of where this term is actually used, we consider as a leak whenever an identifying term is predicted in the right place in the original text or in another place. We refer to the metric named Privacy when measuring the ability of a model to predict any identifying term (i.e., direct or indirect). The metrics named Direct Privacy and Indirect Privacy measure, respectively, the ability of the model to predict only directly or indirectly identifying terms.
(a) N2c2
(b) Hcl
Figure 5. Cumulative distribution of the number of identifiers (both direct and indirect ones), and words per patient: for N2c2 dataset, half of the patients have more than 20 indirect identifiers, and almost all patients have at least 3 indirect identifiers.
@@ 5.4. Comparative Baselines
  • Masked Language Modeling (MLM): First, we consider a classical specialization of a BERT model (named mlm-bert). Next, we consider a baseline that performs the same specialization but using a pseudonymized dataset for the specialization (baseline named mlmA-bert). Finally, we consider two variants of our solution where only direct or indirect identifiers are protected from memorization (baselines named DPPmlm-bert and IPPmlm-bert, respectively).
  • Causal Language Modeling (CLM): We use identical baselines to evaluate our solution. Specifically, we consider a classical specialization of a GPT model (named clm-gpt) that consists in iteratively predicting each term of a sequence of the dataset used for the specialization. We also consider a specialization using a pseudonymized dataset (baseline named clmA-gpt), and two variants of our solution that protect only directly or indirectly identifying terms (baselines named DPPclm-gpt and IPPclm-gptert, respectively).
(a) N2c2
(b) Hcl
@@
5.5. Methodology
This section provides the methodological details we followed to perform our evaluations. First, when datasets are pseudonymized before being used for model finetuning (i.e., in the case of mlmA-bert and clmA-gpt), directly identifying words are replaced by the term ”X” in the text. Then, BERT-like models were trained by randomly replacing a subset of tokens from the sequence by the ¡mask¿ token, and asking the model to predict them using cross-entropy loss. In our setting, 15% of the words are randomly selected. In the case of our solution, identifying words are excluded from this random selection. If a word chosen to be masked consists of several tokens, all of them are masked. At each epoch, 15% of non-identifying words are masked.
Figure 7. Characterization of indirectly identifying words of N2c2 dataset by classifying them into different categories using biomedical named entity recognition (NER): these indirectly identifying terms are present in many categories.
Figure 8. Membership inference capability on patients in the N2c2 dataset by an adversary who has auxiliary information on only two categories (e.g., information on the organs and cancers of its targets).
(a) N2C2
(b) HCL
@@ 5.5. Methodology
All measurements were performed after 4, 8, 16, 32 and 64 epochs to observe their evolution. We used 16 batches of 512 tokens for training. For language modeling, the learning rate starts at 1e-4 and decreases linearly until the end of finetuning. For classification, we took the finetuned model after 64 epochs and finetuned it again by freezing all the BERT layers (to not bias the model and introduce other leakage) except the last one and the classification layer following the methodology described in this study (Sun et al., 2020). The learning rate this time starts at 0, 10% of the total number of warmup steps with the learning rate increasing until reaching 2e-5 then decreasing linearly on the remaining 90% steps.
6. Evaluation
This section presents a comprehensive evaluation of our privacy-preserving language modeling scheme. We first quantify and illustrate the risk of membership inference through the memorization of indirect identifiers (Section 6.1) before evaluating both the utility and privacy trade-off of masked and causal language modeling (Section 6.2 and Section 6.3, respectively). We show that memorization of identifiers can lead to the re-identification of individuals whose data were used in training. Furthermore, we show that protecting the direct identifier only by pseudonymization is not sufficient to protect the model against a risk of memorizing indirect identifiers leading to membership inference, and that it is necessary to protect indirect identifiers as well.
6.1. Illustrating the risk
In this section, we illustrate the risk of membership inference related to the model’s memorization of personal information. More specifically, we analyze the impact of indirectly identifying information in this membership capability. First, we analyze the N2c2 and the Hcl datasets, and analyze the distribution of direct and indirect identifiers, as well as the number of words in each patient’s reports. Figure 5 reports the Cumulative Distribution Function (CDF) of these information. First, we can see that almost all patients have indirect identifiers and half of them have more than twenty indirect identifiers for the N2c2 dataset, and more than ten indirect identifiers for the Hcl dataset. Second, we can see that direct identifiers are about twice as numerous as indirect identifiers. Finally, some indirect identifiers (i.e., words exploited in the documents of a single patient) are also direct identifiers.
Figure 6, in turn, shows for both datasets the cumulative distribution of the number of patients who share words in their reports. The plot distinguishes distinct words from the total number of words by including repetitions. The distribution shows for both datasets that 60% of distinct words in the corpus are used by only one patient and therefore represent indirect identifiers. However, the use of these words represents less than 3% of the total occurrences of words present in all reports. In other words, the 40% of words used in reports of multiple patients represents more than 97% of the total volume of words used.
We now characterize the indirectly identifying words by classifying them into different categories using biomedical named entity recognition (NER) that also includes classes of diseases, chemicals, and genetic entities [^3] https://github.com/librairy/bio-ner. Figure 7 represents for the N2c2 dataset the distribution of indirectly identifying words in these medical-related categories. The distribution shows that indirectly identifying words are classified into a large number of categories where chemicals are the most represented in the N2c2 dataset.
To assess an attacker’s ability to identify that an individual’s data has been used in the specialization of a model from indirect identifier words, we only consider certain categories as plausible auxiliary information. Indeed, some information is less likely to be known by an attacker. Here, we only consider words in the Organ and Cancer categories as words that an adversary could know a priori about the person she is trying to attack. Figure 8 reports the proportion of patients that can be identified from indirectly identifying words in the dataset that fall into the Organ and Cancer categories. Results show that around 15% of the patients can be identified from words in the Cancer category, and 6% from words in the organ category. This identification rate is an upper bound of the risk in the case where the attacker manages to identify that an indirectly identifying word that is part of these two categories of plausible auxiliary information has been learned by a model. In practice, all the words in these two categories may not be known to an attacker. However, thanks to our privacy-preserving language modeling scheme avoiding that an indirectly identifying word cannot be predicted by the model after its specialization, this risk of membership inference is drastically mitigated.
6.2. Utility and Privacy Tradeoff for Masked Language Modeling
We now evaluate the utility and privacy tradeoff for our masked language modeling scheme. Figure 9 illustrates this tradeoff for PPmlm-bert and other comparative baselines for the N2c2 and Hcl datasets. First, the results show that without a specific mechanism to prevent the model from memorizing identifiers (i.e., classical mlm-bert), privacy is significantly degraded, from 0.87 after 4 epochs to 0.66 after 64 epochs. This decrease in privacy means that the model learns more and more identifiers over the epochs and is led to predict them more along the training. Second, training without memorizing indirect identifiers (i.e., IPPmlm-bert) provides about the same level of utility while increasing privacy. This small increase in privacy is due to the reduced number of indirect identifiers. Training without memorizing the direct identifiers, in turn, significantly improves privacy. Indeed, since the number of direct identifiers is significantly larger than the number of indirect identifiers, the increase in privacy is significantly more pronounced. The utility, on the other hand, is similar. Then, training using a pseudonymized dataset (i.e., leveraging a training dataset without direct identifiers, mlmA-bert) provides a similar level of privacy but reduces the utility. This reduction in utility comes from a context used for prediction that does not benefit from direct identifying terms. Finally, our PPmlm-bert solution offers the best utility and privacy tradeoff where the privacy remains stable around 0.98 regardless of the number of epochs, meaning that the model does not learn any identifier directly or indirectly during training. Regarding utility, PPmlm-bert also offers the best utility. This high utility is due to the wealth of information (i.e., using all terms including identifiers) in the context used to predict a term.
(a) Direct Privacy
(b) Indirect Privacy
Figure 10. Evaluating the utility and privacy tradeoff for MLM by biasing the privacy metric toward direct identifiers or indirect identifiers for the N2c2 dataset.
Let us then analyze the behavior of PPmlm-bert and the other baselines if we consider only the prediction of direct identifiers or indirect identifiers as privacy leakage (Figure 10 for the N2c2 dataset). In the former case (Figure 10(a)), results show that mlm-bert and IPPmlm-bert exhibit a close utility and privacy tradeoff. Indeed, since indirect identifiers are not taken into account by the privacy measure, our mechanism for not memorizing indirect identifiers has no observable effect. Consequently, only the solutions that protect direct identifiers exhibit high privacy (i.e., PPmlm-bert, mlmA-bert, and DPPmlm-bert). However, we observe that mlmA-bert exhibits the lowest utility. This lower utility comes from the fact that all direct identifiers are pseudonymized in the training data and are not available in the context to improve the model’s prediction. Once again, PPmlm-bert offers the best trade-off between utility and confidentiality because it benefits from direct identifiers in the context to improve predictions. By comparing the trade-off between mlmA-bert and DPPmlm-bert, we can evaluate the impact of memorization related to the use of directly identifying words in the context for mask prediction. We can see that the privacy impact is very small, and that using a context that exploits direct identifiers slightly improves the utility.
Similarly, if the privacy measure only captures the prediction of indirect identifiers as a leakage (Figure 10(b)), only solutions that carefully protect this category of identifiers exhibit high privacy, this is the case for PPmlm-bert and IPPmlm-bert. The privacy degradation as a function of the number of epochs for the other baselines highlights the fact that they progressively learn indirect identifiers and predict them more and more often during training.
@@ 6.2. Utility and Privacy Tradeoff for Masked Language Modeling
Figure 11. Utility assessment for MLM through a downstream classification task on the N2c2 dataset: all baselines provide roughly the same level of prediction.
@@
We now evaluate the utility of the models through a classification task on the N2c2 dataset. As described in Section 5.3, this classification task aims to predict the patient’s smoking, and the classification models are trained as a downstream task from the specialized model on the dataset. Figure 11 illustrates the prediction accuracy of a classification model after 64 epochs. Counterintuitively, the results show that models that protect identifiers display slightly better accuracy. These results tend to show that identifiers (i.e., information that only affects a single individual) reduce the model’s ability to generalize knowledge and classify correctly.

### v2

@@ 4. Privacy-Preserving Language Modeling
In order to prevent the model from memorizing personal information, the privacy-preserving language modeling that we propose ensures that the choice of mask used during the specialization of a pre-trained model avoids directly and indirectly identifying words. To achieve that, this solution is based on a preprocessing that is responsible for identifying both the direct and indirect identifiers. We first present this preprocessing step (Section 4.1) before explaining how the proposed masked and the causal language modeling work (Section 4.2 and Section 4.3, respectively).
@@ 4.1. Preprocessing: Building a Blacklist
The preprocessing step focuses on identifying identifiers that the model must not to memorize. First, we use an off-the-shelf NER model to directly identify identifying words associated with certain categories. This NER model can be adapted to include more or less categories depending on the field of specialization. Then, we build the bipartite graph $G$ between the individuals of the corpus and the words used in their documents in order to identify indirect identifiers (similar to Figure 3). To do this, once the graph is built, we identify the words that are only linked to one individual. Finally, all these (directly or indirectly) identifiers are kept in a kind of blacklist that is used during the following masked or causal language modeling.
If we make the link with the notion of k-anonymity(Sweeney, 2002), which defines that a data is anonymous if it is used by at least k different users, all the words used as masks to specialize the model in the language modeling step are at least used by two people (i.e., k=2). This makes the task of an adversary wishing to infer the membership of an individual more difficult. We can easily change the exploitation of the bipartite graph to consider a larger value of k for the identification of indirectly identifying information (e.g., words that are linked to 1 or 2 users to have a k=3).
In addition to being able to vary the value of k to protect indirect identifiers, and to use a NER adapted to the domain of specialization to identify direct identifiers, it should be noted that the cost of this preprocessing is reduced.
@@ 4.2. Privacy-Preserving Masked Language Modeling
Compared to the standard masked language modeling scheme that randomly chooses terms to mask during supervised learning, our privacy-preserving masked language modeling ensures that directly or indirectly identifying terms are never masked by leveraging the blacklist constructed during the preprocessing step. More precisely, the choice of the mask is made randomly among the words that are not in the blacklist.
@@ 4.3. Privacy-Preserving Causal Language Modeling
Similar to privacy-preserving masked language modeling, our privacy preserving causal language modeling ensures that masked words are not directly or indirectly identifying words. However, since masked words are always at the end of the sequence to be predicted, we exploit the blacklist created during the preprocessing step to check that this last term is not an identifier. If it is, we modify it and replace it with a padding token and add the next word for prediction. The advantage of using a padding token is that it is not used in the loss calculation during the fine-tuning, and therefore does not degrade the model’s predictions.
@@ 5. Experimental Setup
We performed an extensive evaluation of our privacy-preserving language modeling scheme using medical datasets (Section 5.1) and considering both off-the-shelf BERT and GPT pre-trained models (Section 5.2). To capture the full impact of our proposed solutions, we considered a set of both utility and privacy metrics (Section 5.3) and considered different comparative baseline approaches (Section 5.4). Finally, we describe our methodology (Section 5.5).
@@ 5.1. Datasets
To conduct our experiments on health data, we considered the N2c2 (National NLP Clinical Challenges) datasets (Henry et al., 2020). More precisely, we leverage two datasets gathering medical discharge summaries (i.e., english free text). In the first one (Uzuner et al., 2007), 928 records have been annotated for replacing all authentic Personally Identifiable Information (PII) with realistic surrogates. These PIIs fall into different categories: six of the seventeen textual PII categories listed by HIPAA (only these six categories appear in the data: patients, locations, dates, IDs, phone numbers, and ages), and two additional categories, doctors and hospitals, resulting in eight PII categories in the dataset. In the second dataset (Uzuner et al., 2008) almost 500 records have been annotated by pulmonologists. The pulmonologists were asked to classify patient records into five possible smoking status categories as past smoker, current smoker, smoker, non-smoker, and unknown. We leverage this dataset to assess the utility degradation of our privacy-preserving language modeling through a downstream classification task.
@@ 5.2. Model Architecture
In order to cover the different use cases of language models (i.e., the generative models and the descriptive text models), we consider both the BERT architecture (i.e., the base model case composed of 12 layers, 768 hidden dimensions, 12 attention heads, 110M parameters) and the GPT-2 architecture (Radford et al., 2019) (the smallest version of GPT-2, with 124M parameters) from Hugging Face777 Hugging Face - https://huggingface.co/. We did not consider larger model versions due to resource limitations, and we also discarded distilled version of these models (Sanh et al., 2020b).
@@ 5.3. Metrics
We consider both utility and privacy measures in our evaluation to better assess the trade-off between protecting training data and the performance of the model in language modeling.
5.3.1. Utility
For the utility evaluation of BERT-like models, we quantify the ability of the models to predict masked terms well (measured by the prediction accuracy), and the model’s performance in classification by creating a downstream task (measured by the F1-Score metric on the prediction of the smoking status). For GPT-like models, the utility is evaluated by the model’s ability to generate the right tokens in the right places (measured by the tokens accuracy). To evaluate the quality of the generated text, we also consider the BLEU and ROUGE scores. While the BLEU (Bilingual Evaluation Understudy) score measures the similarity between the generated text and the original text using n-grams (i.e., contiguous sequences of n words), the ROUGE (Recall-Oriented Understudy for Gisting Evaluation) score calculates the recall of common n-grams.
5.3.2. Privacy
To assess the model’ privacy, we study its ability to predict identifiers (direct and indirect) contained in the dataset used to specialize the model. An individual’s presence is revealed as soon as at least one of their identifiers is predicted or regurgitated by the model, whether this prediction is in the correct place in the original text or elsewhere. We use the metric Privacy to measure the ability of a model to predict any identifying term (i.e., direct or indirect identifier). A Privacy=1 means that no identifiers are predicted or regurgitated by the model, and a Privacy=0 means that all identifiers are leaked. The metrics named Direct Privacy and Indirect Privacy measure, respectively, the ability of the model to predict only directly or indirectly identifying terms. We also considered MIAs to measure an adversary’s ability to correctly predict whether a patient’s reports were used in model fine-tuning. Specifically, we considered a state-of-the-art MIA attack based on the loss (Duan et al., 2024) through the MIMIR package [^8] MIMIR - https://github.com/iamgroot42/mimir, and an attack that infers membership if the model predicts any identifiers present in the adversary’s auxiliary knowledge.
@@ 5.4. Comparative Baselines
  • Masked Language Modeling (MLM): First, we consider a classical specialization of a BERT model (named mlm-bert). Next, we consider a baseline that performs the same specialization but using a pseudonymized dataset for the fine-tuning (baseline named mlmA-bert). Finally, we consider two variants of our solution where only direct or indirect identifiers are protected from memorization (baselines named DPPmlm-bert and IPPmlm-bert, respectively).
  • Causal Language Modeling (CLM): We consider a classical specialization of a GPT model (named clm-gpt) that consists in iteratively predicting each term of a sequence of the dataset used for the specialization. We also consider a specialization using a pseudonymized dataset (baseline named clmA-gpt), and two variants of our solution that protect only directly or indirectly identifying terms (baselines named DPPclm-gpt and IPPclm-gpt, respectively). Lastly, we consider a state-of-the-art baseline using a differentially private fine-tuning of pre-trained language models (Yu et al., 2022; Wutschitz et al., 2022) integrating the Opacus library [^9] Opacus - https://opacus.ai/.
5.5. Methodology
This section provides the methodological details we followed to perform our evaluations. First, when datasets are pseudonymized before being used for model fine-tuning (i.e., in the case of mlmA-bert and clmA-gpt), direct identifiers are replaced by the term ”X” in the text. Then, BERT-like models were trained by randomly replacing a subset of tokens from the sequence by the ¡mask¿ token, and asking the model to predict them using cross-entropy loss. In our setting, 15% of the words are randomly selected. In the case of our solution, identifying words are excluded from this random selection. If a word chosen to be masked consists of several tokens, all of them are masked. At each epoch, 15% of non-identifying words are masked.
All measurements were performed after 4, 8, 16, 32 and 64 epochs to observe their evolution. We used 16 batches of 512 tokens for training. For language modeling, the learning rate starts at 1e-4 and decreases linearly until the end of finetuning. For classification, we took the finetuned model after 64 epochs and finetuned it again by freezing all the BERT layers (to not bias the model and introduce other leakage) except the last one and the classification layer following the methodology described in this study (Sun et al., 2020). The learning rate this time starts at 0, 10% of the total number of warmup steps with the learning rate increasing until reaching 2e-5 then decreasing linearly on the remaining 90% steps. All the computation has been parallelized on a hybrid GPU/CPU computing farm.
Figure 5. Cumulative distribution of the number of identifiers (both direct and indirect ones), and words per patient: half of the patients have more than 20 indirect identifiers, and almost all patient have at least 3 indirect identifiers.
6. Evaluation
This section presents a comprehensive evaluation of our privacy-preserving language modeling scheme. We first analyze the distribution of identifiers in the dataset (Section 6.1). We then evaluate the utility and privacy trade-off of both masked and causal language modeling (Section 6.2 and Section 6.3, respectively). Finally, we evaluate the impact of MIAs (Section 6.4) before evaluating the impact of considering n-grams as potential indirect identifiers (Section 6.5). We show that memorization of identifiers can lead to the membership inference of individuals whose data were used in training. Furthermore, we show that protecting the direct identifier only through pseudonymization is not enough to protect the model against a risk of memorizing indirect identifiers.
6.1. Identifiers Analysis
In this section, we first analyze the distribution of direct and indirect identifiers in the N2c2 dataset, as well as the number of words in each patient’s reports (Figure 5). We can see that almost all patients have indirect identifiers and half of them have more than twenty indirect identifiers. We can also see that direct identifiers are about twice as numerous as indirect identifiers. Finally, a fraction of indirect identifiers (i.e., words exploited in the documents of a single patient) are also direct identifiers.
@@ 6.1. Identifiers Analysis
Figure 6, in turn, shows the cumulative distribution of the number of patients who share words in their reports. The plot distinguishes distinct words from the total number of words by including repetitions. The distribution shows that 60% of distinct words in the corpus are used by only one patient and therefore represent indirect identifiers. However, the use of these words represents less than 3% of the total occurrences of words present in all reports. In other words, the 40% of words used in reports of multiple patients represents more than 97% of the total volume of words used.
6.2. Utility and Privacy Tradeoff for Masked Language Modeling
@@
Figure 8. Utility assessment for MLM through a down-stream classification task: all baselines provide roughly the same level of F1-Score.
We now evaluate the utility and privacy tradeoff for our masked language modeling scheme. Figure 7 illustrates this tradeoff for PPmlm-bert and other comparative baselines. First, the results show that without a specific mechanism to prevent the model from memorizing identifiers (i.e., classical mlm-bert), privacy is significantly degraded, from 0.87 after 4 epochs to 0.66 after 64 epochs. This decrease in privacy means that the model learns more and more identifiers over the epochs and is led to predict them more along the training. Second, training without memorizing indirect identifiers (i.e., IPPmlm-bert) provides about the same level of utility while increasing privacy. This small increase in privacy is due to the reduced number of indirect identifiers. Training without memorizing the direct identifiers, in turn, significantly improves privacy. Indeed, since the number of direct identifiers is significantly larger than the number of indirect identifiers, the increase in privacy is significantly more pronounced. The utility, on the other hand, is similar. Then, training using a pseudonymized dataset (i.e., leveraging a training dataset without direct identifiers, mlmA-bert) provides a similar level of privacy but reduces the utility. This reduction in utility comes from a context used for prediction that does not benefit from direct identifying terms. Finally, our PPmlm-bert solution offers the best utility and privacy tradeoff where the privacy remains stable around 0.98 regardless of the number of epochs, meaning that the model does not learn any identifier directly or indirectly during the fine-tuning. Regarding utility, PPmlm-bert also offers the best utility. This high utility is due to the wealth of information (i.e., using all terms including identifiers) in the context used to predict a term. In Appendix A.1, we examine the behavior of PPmlm-bert and other baselines if privacy leakages are measured only from direct or indirect identifiers.
@@
We now evaluate the utility of the models through a classification task. As described in Section 5.3, this classification task aims to predict the patient’s smoking, and the classification models are trained as a downstream task from the specialized model on the dataset. Figure 8 illustrates the F1-Score of the classification model after 8 epochs. Counterintuitively, the results show that models that protect identifiers display slightly better accuracy. These results tend to show that identifiers (i.e., information that only affects a single individual) reduce the model’s ability to generalize knowledge and classify correctly.